OpenStudy (anonymous):

the perimeter of a rectangle is twice the sum of its length and its width. the perimeter is 46 meters and its length is 2 meters more than twice its width. What is the length?

OpenStudy (ajprincess):

Let the perimeter of the rectangle be P, length be l and width be w. l=2+2w P=2(l+w) 46=2(2+2w+w) Can u find w @elexusvanderhorst?

OpenStudy (anonymous):

w=7?

OpenStudy (ajprincess):

yup:)

OpenStudy (anonymous):

so how would i turn this into an equation?

OpenStudy (ajprincess):

turn vat?

OpenStudy (anonymous):

oh wait sorry so my answer would be w=7?

OpenStudy (ajprincess):

no u r askd to find l. plug in the value of w in l=2+2w and find l:)

OpenStudy (anonymous):

oh ok thank you so much

OpenStudy (ajprincess):

Welcome:) If u dnt get any part of it pls ask:)